# Cross-Modal and Uncertainty-Aware Agglomeration for Open-Vocabulary 3D Scene Understanding

### v1

@@ 4 Experiments
We run extensive experiments over a wide set of tasks, going from open-vocabulary segmentation and cross-domain generalization to the evaluation with fine-grained class vocabularies. This section is organized as follows. Sec. 4.14.2 provide the dataset and implementation details of  CUA-O3D used in our experiments. Sec. 4.3 and Sec. 4.4 illustrate the potential when embracing various 2D foundation models and evaluate the open-vocabulary 3D semantic segmentation, compared with related methods. Sec. 4.5 reports cross-domain generalization under common and fine-grained class evaluation while sec 4.6 presents the downstream performance after distillation. The final Sec. 4.7 analyzes the improvements of each proposed components.
@@ 4.2 Implementation Details
We implement our method in the PyTorch framework [63] and employ experiments on a single NVIDIA A100 GPU for ScanNetV2 and Matterport3D, respectively. We follow  [64] and use MinkUNet18A [13] as our 3D backbone starting from randomly initialized weights and LSeg [6] as our pre-trained VLM. During our training, we adopt Adam optimizer [39] with an initial learning rate of $41e-$ 4 and an exponential decay and train our pipeline for 50 epochs. For the teacher model parameter update, we set the momentum coefficient $\beta$ to $0.99$ and $\gamma$ to $1$. Besides, we employ the voxel size of $2$cm and batch size of $2$ for both the ScanNetV2 and Matterport3D experiments. Due to the GPU memory limitation, we uniformly sample $20k$ point features to train the model and input only the 3D point position without RGB information to the MinkowskiNet. We utilize random horizontal flip and elastic distortion as data augmentations over point clouds while following BP-Net [32] to apply color, jitter, and hue feature transformations over 2D feature embeddings.
@@ 4.3 Qualitative Comparisons
To demonstrate the motivation of various 2D foundation models agglomeration, we investigate the results using K-Means to cluster the projected feature embeddings from different 2D models and our distilled features as shown in Fig. 4. As can be observed that different 2D model performs heterogeneous yet complementary results. Likewise, the table from the top-left sample in Fig. 4 displays various clustering results, leading to our agglomerated model being able to output more accurate and consistent clusters. Meanwhile, we also utilize UMAP [72] to better illustrate the intrinsic characteristics when adopting a specific 2D model to employ the 3D model distillation, as the output feature embeddings from various dimensions of DINOv2 and Stable Diffusion are not elaborated for direct matching with text embeddings. As shown on the right side of Fig. 4, DINOv2 indicates smoother and more consistent results though Lseg has been trained to align with the text encoder within dense supervision. Stable Difusion presents intriguing geometric characteristics, all of which are capable of agglomerating potential foundation 3D models using distillation.
@@ 4.4 Open-Vocabulary 3D Semantic Segmentation
To showcase that our method CUA-O3D can boost the performances of the open-vocabulary 3D semantic segmentation model, we compare our method with existing works within two types of settings, including fully supervised (Fully-sup.) and zero-shot. From Table 2, our method surpasses the recent work OpenScenereimple [64] with $+2.5\%$ mIoU under 3D-distill and $+3.1\%$ mIoU under 2D3D-ensemble on ScanNetV2 val set, and $+0.8\%$ mIoU under 3D-distill sand $+0.7\%$ mIoU under 2D3D-ensemble on Matterport3D val set, respectively. This further proves our method not only distills heterogeneous yet complementary knowledge but also reconciles with noisy supervision. This is achieved by the proposed deterministic uncertainty estimation which adaptively captures inherent noise and then weights the corresponding distillation.
@@
Supervised by various 2D foundation models, like Lseg, DINOv2, and Stable Diffusion, the 3D model learns to align with the open-vocabulary feature space together with spatial and geometric awareness. Some open-vocabulary semantic segmentation visualizations are shown in Fig. 5. Additional experiments with AMRADIO [67] can be referred to in our supplementary material.
Figure 5: Open-vocabulary semantic segmentation comparisons in terms of ScanNetV2 and Matterport. Our approach displays superior performance over the OpenScene, which is regarded as our baseline. Best view zoom in and out.
@@
4.5 Cross-Dataset Generalization
We study the generalization capability of our CUA-O3D to unseen datasets without further fine-tuning (i.e., zero-shot). This is achieved by evaluating the cross-dataset performance when training and evaluating different datasets. We use ScanNetV2 and Matterport3D as training and evaluation datasets and analyze the cross-dataset results when training on ScanNetV2 and evaluating on Matterport3D, and vice versa. Table 3 and Table 4 report the cross-dataset results and with different granularities, respectively.
@@ 4.5 Cross-Dataset Generalization
Table 5: Experimental results on ScanNetV2 and Matterport in terms of val on linear probing evaluation. Upperbound-full sup. denotes the fully-supervised upperbounding results while Baseline init. means initialize the model from our baseline model and then perform linear probing evaluation.
@@ 4.6 Linear Probing
In this section, we exploit how the trained 3D model will perform after agglomerating various 2D models. We then conduct experiments that employ linear probe learning based on the 3D model after distilling from various 2D models. Specifically, we construct a simple MLP layer on top of the frozen 3D model backbone and train the linear layer only following the fully-supervised manner. As shown in Table 5, the method concatenates all three mapping layers corresponding to Lseg, DINOv2, and Stable Diffusion, and then maps the concatenated features to close set label spaces, we can see this way can achieve the best performances, 62.1% mIoU and 45.8% mIoU on ScanNetV2 and Matterport val after tuned on train. Note that this obtains 7.7% mIoU improvement over the model initialized from our baseline model. We can also observe that simply mapping the DINOv2 layer from the 3D model attains very competitive segmentation performance while mapping the Lseg layer only which has been trained before to align with the text encoder realizes an inferior one. These further insights that we shall seek more suitable 2D model selections to help develop potential foundational 3D models, whereas DINOv2 presents strong generalizability and flexibility, which is consistent with the observation [55, 20].
@@ 4.7 Ablative Studies
In this section, we study the improvements from each proposed component. As shown in Table 6, we begin by gradually adding each component to the 3D model training and find that only combining with Lseg and DINOv2 leads to marginal open-vocabulary semantic segmentation which can be conjectured that DINOv2 has not been aligned with language space before though it excels at spatial perception ability. Then, the performance is boosted by 1.3% mIoU when introducing Stable Diffusion supervision, while further improved by 2.1% mIoU and 1.9% mAcc via our proposed deterministic estimation to help the model adaptively harmonize the heterogeneous knowledge from various 2D models. Interestingly, if we apply auto-weighting to enable the model to learn by itself, the model training falls into collapse, which we surmise is due to the minimal optimization objective and the model quickly gets into a trivial solution. Overall, CUA-O3D can improve the 3D model only from 51.4% mIoU and 62.3% mAcc to 54.1% mIoU and 64.1% mAcc, which further demonstrates the effectiveness of our method.
@@ 5 Conclusions
In this paper, we first investigate the cross-modal agglomeration from various 2D foundation models into one 3D model, in pursuit of a potential foundational 3D model. To resolve the heterogeneous bias and inherent noise from 2D feature supervisions, we then propose a deterministic uncertainty estimation to capture model-specific uncertainties across diverse semantic and geometric sensitivities, which is then leveraged to weight the corresponding distillation loss adaptively. In this way, the trained 3D model performs competitive open-vocabulary segmentation while achieving robust cross-domain alignment and strong spatial perception ability, which hopes to shed new light on the community.
Acknowledgments
This work was supported by the MUR PNRR project FAIR (PE00000013) funded by the NextGenerationEU and the EU Horizon project ELIAS (No. 101120237). We acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support.
@@ 6 Distillation Loss Choice
Table 10: Experimental results on ScanNetV2 and Matterport in terms of val on linear probing evaluation. Upperbound-full sup. denotes the fully-supervised upperbounding results while Baseline init. means initialize the model from our baseline model and then perform linear probing evaluation. Adding AMRADIO for comparision.
@@ 10 More Visualizations
In this section, we demonstrate more clustering results for ScanNetV2 and Matterport to further clarify the motivation of agglomerating heterogeneous and complementary feature supervisions from various 2D foundation models in Fig. 6 and Fig. 7, while open-vocabulary semantic segmentation visualizations are displayed in Fig. 9 and Fig. 10.
@@ 11 Limitations and Future Improvements
Figure 7: Clustering visualizations of Matterport from various 2D foundation models.
@@
Figure 10: Open-vocabulary semantic segmentation visualizations of Matterport.

### v2

@@ 4 Experiments
We run extensive experiments over a wide set of tasks, going from open-vocabulary segmentation and cross-domain generalization to the evaluation with fine-grained class vocabularies. This section is organized as follows. Sec. 4.14.2 provide the dataset and implementation details of  CUA-O3D used in our experiments. Sec. 4.3 and Sec. 4.4 illustrate the potential when embracing various 2D foundation models and evaluate the open-vocabulary 3D semantic segmentation, compared with related methods. Sec. 4.5 reports cross-domain generalization under common and fine-grained class evaluation while sec 4.6 presents the downstream performance after distillation. The final Sec. 4.7 ablates and analyzes the improvements of each proposed components.
@@ 4.2 Implementation Details
We implement our method in the PyTorch framework [63] and employ experiments on a single NVIDIA A100 GPU for ScanNetV2 and Matterport3D, respectively. We follow  [64] and use MinkUNet18A [11] as our 3D backbone starting from randomly initialized weights, and Lseg [5] as our pre-trained VLM. During our training, we adopt Adam optimizer [37] with an initial learning rate of $41e-$ 4 and an exponential decay to train our pipeline for 50 epochs. For the teacher model parameter update, we set the momentum coefficient $\beta$ to $0.99$ and $\gamma$ to $1$. Besides, we employ the voxel size of $2$cm and batch size of $2$ for both the ScanNetV2 and Matterport3D experiments. Due to the GPU memory limitation, we uniformly sample $20k$ point features to train the model and input only the 3D point position without RGB information to the MinkowskiNet. We utilize random horizontal flip and elastic distortion as data augmentations over point clouds while following BP-Net [30] to apply color, jitter, and hue feature transformations over 2D feature embeddings.
@@ 4.3 Qualitative Comparisons
To demonstrate the motivation of various 2D foundation models agglomeration, we investigate the results using K-Means to cluster the projected feature embeddings from different 2D models and our distilled features as shown in Fig. 4. As can be observed that different 2D model performs heterogeneous yet complementary results. Likewise, the table from the top-left sample in Fig. 4 displays various clustering results, leading to our agglomerated model being able to output more accurate and consistent clusters. Meanwhile, we also utilize UMAP [72] to better illustrate the intrinsic characteristics when adopting a specific 2D model to employ the 3D model distillation, since the output feature embeddings from various dimensions of DINOv2 and Stable Diffusion are not elaborated for direct matching with text embeddings. As shown on the right side of Fig. 4, DINOv2 indicates smoother and more consistent results though Lseg has been trained to align with the text encoder before within dense supervision. Stable Difusion presents intriguing geometric characteristics, all of which are capable of agglomerating potential foundation 3D models using distillation.
@@ 4.4 Open-Vocabulary 3D Semantic Segmentation
To showcase that our method CUA-O3D can boost the performances of the open-vocabulary 3D semantic segmentation model, we compare our method with existing works within two types of settings, including fully supervised (Fully-sup.) and zero-shot. From Table 2, our method surpasses the recent work OpenScenereimple [64] with $+2.5\%$ mIoU under 3D-distill and $+3.1\%$ mIoU under 2D3D-ensemble on ScanNetV2 val set, and $+0.8\%$ mIoU under 3D-distill and $+0.7\%$ mIoU under 2D3D-ensemble on Matterport3D val set, respectively. This further proves our method not only distills heterogeneous yet complementary knowledge into the 3D model but also reconciles with 2D noisy supervisions. This is achieved by the proposed deterministic uncertainty estimation which adaptively captures inherent noise and then weights the corresponding distillation. Supervised by various 2D foundation models, like Lseg, DINOv2, and Stable Diffusion, the 3D model learns to align with the open-vocabulary features together with spatial and geometric awareness. Some open-vocabulary 3D semantic segmentation visualizations are shown in Fig. 5. Additional experiments with AMRADIO [67] can be referred to our supplementary material.
@@
4.5 Cross-Dataset Generalization
We study the generalization capability of our CUA-O3D to unseen datasets without further fine-tuning (i.e., zero-shot). This is achieved by evaluating the cross-dataset performance when training and evaluating different datasets. We use ScanNetV2 and Matterport3D as training and evaluation datasets respectively, and analyze the cross-dataset results when training on ScanNetV2 and evaluating on Matterport3D, and vice versa. Table 3 and Table 4 report the cross-dataset results and with different granularities, respectively.
Figure 5: Open-vocabulary 3D semantic segmentation comparisons in terms of ScanNetV2 and Matterport3D. Our approach displays superior performance over the OpenScene, which is regarded as our baseline. Best view zoom in and out.
@@ 4.5 Cross-Dataset Generalization
Table 5: Experimental results on ScanNetV2 and Matterport3D in terms of val on linear probing evaluation. Upperbound-full sup. denotes the fully-supervised upperbounding results while Baseline init. means initialize the model from our baseline model and then perform linear probing evaluation.
@@ 4.6 Linear Probing
In this section, we exploit how the trained 3D model will perform after agglomerating various 2D models. We then conduct experiments that employ linear probe learning based on the 3D model after distilling from various 2D models. Specifically, we construct a simple MLP layer on top of the frozen 3D model backbone and train the linear layer only following the fully-supervised manner. As shown in Table 5, the method concatenates all three mapping layers corresponding to Lseg, DINOv2, and Stable Diffusion, and then maps the concatenated features to close set label spaces, we can see this way can achieve the best performances, 62.1% mIoU and 45.8% mIoU on ScanNetV2 and Matterport3D val after tuned on train. Note that this obtains 7.7% mIoU improvement over the model initialized from our baseline model. We can also observe that simply mapping the DINOv2 layer from the 3D model attains very competitive segmentation performance while mapping the Lseg layer only which has been trained before to align with the text encoder realizes an inferior one. These further insights that we shall seek more suitable 2D model selections to help develop potential foundational 3D models, whereas DINOv2 presents strong generalizability and flexibility, which is consistent with the observation [53, 18].
@@ 4.7 Ablative Studies
In this section, we study the improvements from each proposed component. As shown in Table 6, we begin by gradually adding each component to the 3D model training and find that only combining with Lseg and DINOv2 leads to marginal open-vocabulary 3D semantic segmentation which can be conjectured that DINOv2 has not been aligned with language space before though it excels at spatial perception ability. Then, the performance is boosted by 1.3% mIoU when introducing Stable Diffusion supervision, while further improved by 2.1% mIoU and 1.9% mAcc via our proposed deterministic estimation to help the model adaptively harmonize the heterogeneous knowledge from various 2D models. Interestingly, if we apply auto-weighting to enable the model to learn by itself, the model training falls into collapse, which we surmise it is due to the minimal optimization objective and the model quickly gets into a trivial solution. Overall, CUA-O3D can improve the 3D model only from 51.4% mIoU and 62.3% mAcc to 54.1% mIoU and 64.1% mAcc, which further demonstrates the effectiveness of our method.
@@ 5 Conclusions
In this paper, we first investigate the cross-modal agglomeration from various 2D foundation models into one 3D model, in pursuit of a potential foundational 3D model. To resolve the heterogeneous bias and inherent noise from 2D feature supervisions, we then propose a deterministic uncertainty estimation to capture 2D model-specific uncertainties across diverse semantic and geometric sensitivities, which is then leveraged to weight the corresponding distillation loss adaptively. In this way, the trained 3D model performs competitive open-vocabulary segmentation while achieving robust cross-domain alignment and strong spatial perception ability, which hopes to shed new light on the community.
Acknowledgments This work was supported by the MUR PNRR project FAIR (PE00000013) funded by the NextGenerationEU and the EU Horizon project ELIAS (No. 101120237). We acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support.
@@ 6 Distillation Loss Choice
Table 10: Experimental results on ScanNetV2 and Matterport3D in terms of val on linear probing evaluation. Upperbound-full sup. denotes the fully-supervised upperbounding results while Baseline init. means initialize the model from our baseline model and then perform linear probing evaluation. Adding AMRADIO for comparision.
@@ 10 More Visualizations
In this section, we demonstrate more clustering results for ScanNetV2 and Matterport3D to further clarify the motivation of agglomerating heterogeneous and complementary feature supervisions from various 2D foundation models in Fig. 6 and Fig. 7, while open-vocabulary semantic segmentation visualizations are displayed in Fig. 9 and Fig. 10.
@@ 11 Limitations and Future Improvements
Figure 7: Clustering visualizations of Matterport3D from various 2D foundation models.
@@
Figure 10: Open-vocabulary semantic segmentation visualizations of Matterport3D.